OpenStudy (anonymous):

The perimeter of a rectangle is 36 m. If its length is increased by 1 m and the width is increased by 2m, its area will increase by 30m^2. Find the area of the original rectangle.

OpenStudy (bahrom7893):

(l+1)*(w+2)=30+a l*w=a 2(l+w) = 36

OpenStudy (bahrom7893):

find a from there

OpenStudy (anonymous):